• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Templon

    Log entry time 13:57:48 on June 2,2000

    Entry number 42880

    This entry is a followup to: 42828

    keyword=vdc problem "solved"

    Talked to Bogdan about VDC plots from last night. He indicated that
    the problem has to do with a threshold too high on e-arm. When control
    was lost, a 'reasonable value' was set locally before Hall was closed.
    The value was reasonable but not optimal. Threshold should be lower.
    Also explains resolution problem (but not why resolution is worse on V1
    than on U1). Should call Bogdan to reduce e-arm threshold at next
    hut access.